OpenStudy (anonymous):

4a+7b+3a-2b+2a ; a = -5, b = 3 I got 5?? Help :c

OpenStudy (jhannybean):

4a +7b +3a -2b +2a = 4(-5) +7(3) +3(-5) -2(3)+2(-5) =

OpenStudy (triciaal):

regroup substitute simplify

OpenStudy (jhannybean):

so just punch this all in to your calculator

OpenStudy (anonymous):

I got -30 , which is one of my answer choices!! yay thank you so much(:

OpenStudy (jhannybean):

No problem :)

OpenStudy (triciaal):

(4 + 3 + 2)*(-5) + (7-2)*3 = 9*-5 + 5*3 = -45 + 15 = -30

OpenStudy (anonymous):

can you help with one more? ill post it and tag you @Jhannybean

OpenStudy (jhannybean):

sure